### v1

# Bridging the Social & Technical Divide in Augmentative and Alternative Communication (AAC) Applications for Autistic Adults

Abstract.
Natural Language Processing (NLP) techniques are being used more frequently to improve high-tech Augmentative and Alternative Communication (AAC), but many of these techniques are integrated without the inclusion of the users’ perspectives. As many of these tools are created with children in mind, autistic adults are often neglected in the design of AAC tools to begin with. We conducted in-depth interviews with 12 autistic adults to find the pain points of current AAC and determine what general technological advances they would find helpful. We found that in addition to technological issues, there are many societal issues as well. We found 9 different categories of themes from our interviews: input options, output options, selecting or adapting AAC for a good fit, when to start or swap AAC, benefits (of use), access (to AAC), stumbling blocks for continued use, social concerns, and lack of control. In this paper, we go through these nine categories in depth and then suggest possible guidelines for the NLP community, AAC application makers, and policy makers to improve AAC use for autistic adults.
accessibility, augmentative and alternative communication, AAC, ASD, autism, NLP, word prediction, large language models, LLMs, ChatGPT
copyright: acmlicensedjournalyear: 2024ccs: Human-centered computing Accessibility technologiesccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Accessibility design and evaluation methodsccs: Human-centered computing Natural language interfacesccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics People with disabilities
1. Introduction
Thanks to the neurodiversity movement online (Kapp, 2020; Botha et al., 2024), many people who were previously overlooked for diagnoses as children are now realizing that they are autistic as adults. Since these autistic individuals are coming into their diagnoses well into adulthood, they are learning about accommodation much later in life. This means that tools like Augmentative and Alternative Communication (or AAC), which replace or supplement a person’s oral communication, are discovered by people who 1) already know how to read and write, 2) have been speaking orally (to their detriment) for years (Zisk and Dalton, 2019), and 3) may not have access to professionals such as speech-language pathologists. Furthermore, the tools themselves are usually developed with children in mind, with speech-language pathologists teaching children how to use AAC to speak their first language.
Even though there has been recent recognition within AAC research showing that AAC use can build agency (Valencia et al., 2020) and garner empathy from the people in the AAC user’s life (van Grunsven and Roeser, 2022), autistic adults are largely ignored within AAC research (Ganz et al., 2012; Lorah et al., 2022; Trembath et al., 2014) and very few are involved in the research of new AAC tools (Bastable et al., 2021). Because of this, autistic adults have learned to navigate a world of AAC tools that are not made for them.
On the technological side of things, research in high-tech AAC tools [^1] AAC tools can range across the spectrum of technology levels—from writing, spelling out words on a letter-board, or using cards with pictures on them; to tablet/phone-based applications that cost several hundreds of dollars. has seen a recent uptick in interest from the Natural Language Processing (NLP) community due to the introduction of large language models (LLMs) like ChatGPT to the general public (Cai et al., 2022; Kumar et al., 2022; Valencia et al., 2023; Yusufali et al., 2023). However, we should tread carefully as NLP and other machine learning techniques can cause harm if not implemented correctly. There are crucial privacy issues such as training on users’ data with out their knowing or sending/saving data to “the cloud” with questionable levels of security. Furthermore, implementing NLP techniques blindly without the inclusion of the community of users can backfire. Thankfully, there has been a push to include AAC users as contributors in AAC development (Bastable et al., 2021; Beneteau, 2020). Our work continues down this path by beginning conversations with AAC users to bridge the gap between what can be done technologically and what users actually want, with a focus on high-tech, computer-based AAC applications.
In this paper, we wanted to learn about the needs of autistic adults such that we can improve AAC applications and dig into some of the deeper issues as to why AAC tools are not always available. In the vein of looking to improve AAC tools for a sub-community of users who are often unheard—autistic adults, we were interested in the following research questions:
  • RQ1: What are pitfalls and challenges that autistic AAC users run into when using existing AAC applications?
  • RQ2: What types of features do autistic adults require or prefer to see from AAC applications?
  • RQ3: How does the person’s identity as an autistic person factor into their AAC use?
  • RQ4: How do participants feel about NLP techniques being integrated into their AAC apps?
To learn more about this intersection of identities, we ran 12 in-depth semi-structured interviews with autistic adults who were recruited online. We outline some themes that we have identified throughout the interviews and provide design recommendations to address them from the perspective of three different societal levels. To match the language autistic adults use (Zisk and Konyn, 2022), we will be using the terms ”speaking orally”, ”using mouth-words”, or ”speaking” (vs ”nonspeaking”) interchangeably throughout the paper to refer to using one’s vocal tract to produce language.
2. Related Work
2.1. Autistic Adult AAC Users
As mentioned in the Introduction, the majority of AAC work either involves autistic children or adults with physical disabilities. However, there is a recent (although small) cohort of work that has recognized this intersectional identify of adult autistic users of AAC (Donaldson et al., 2021; Hajjar and McCarthy, 2022; Holyfield et al., 2020; Hughes et al., 2022; Zisk and Dalton, 2019). Even so, a lot of this work has a limited context. For example, Holyfield et al. (2017) performed a literature review of introductions of AAC (“interventions”) into the lives of autistic adolescents and adults. Out of the 19 participants mentioned in the papers they found, only 4 of them were adults.
Other works have focused on adults with intellectual and developmental disabilities broadly (Holyfield et al., 2020; Hughes et al., 2022), which includes people with autism spectrum disorder. Previous work includes using AAC to improve users’ skills such as single word sight reading (Holyfield et al., 2020) or achieving their own personalized communication goals (Hughes et al., 2022). Hajjar and McCarthy (2022) asked adult AAC users how they experienced using AAC during physical activities like sports.
To our knowledge, there have only been two papers written so far looking specifically at autistic adults’ general AAC use (Zisk and Dalton, 2019; Donaldson et al., 2021)—both papers focused on speaking adults. Our study mirrors these works since we looked at autistic adults who were not “completely nonspeaking”. (However, we speculate that it is probably difficult for someone to consider themselves fully nonspeaking if they have forced speech for most of their life and only encountered AAC as an adult, as many of our participants have.) Zisk and Dalton (2019) reviewed academic and non-academic sources to find recurring themes of how speaking autistic adult AAC users communicate when they can occasionally use speech, what types of tools they use to communicate (both specialized for AAC and not), and any barriers to using AAC. Donaldson et al. (2021) performed an open-ended written survey with six participants, asking questions specific to AAC use history, access, attitudes to use, and any recommendations they might have for people who might use or interact with people who use AAC. While our work has rediscovered some of the same social themes (e.g., choice of AAC, autonomy, acceptance), we also dive into technical and privacy themes.
2.2. Technological Advances in AAC
Technologically speaking, most research on improving high-tech AAC tools focuses on customizablity or automating/predicting what the user means to do (Curtis et al., 2022). Much of the prediction involves predicting the text that the user would want to say. In the field of natural language processing (NLP), this primarily involves the use of a “language model”. Language modeling is a way of creating a probabilistic model of text. That is, after training on many documents, they can learn to predict what word(s) will come next given the words it has seen so far. For AAC, research has focused on next word prediction (Adhikary et al., 2019; Freedman et al., 2013; Schadle, 2004; Trnka et al., 2007; Wiegand and Patel, 2012), contextual or discourse-based prediction (Freedman et al., 2013; Mitchell and Sproat, 2012; Perkoff, 2021), personalized language (Di Mascio et al., 2019; Freedman et al., 2013; Li et al., 2022; Valencia et al., 2023), or generating sentences from a set of keywords (Cai et al., 2022; Kumar et al., 2022). Other unique approaches include using automatic speech recognition (speech-to-text) to improve word prediction by listening to AAC users’ conversation partners and using the recognized words as additional context (Adhikary et al., 2019); and classifying emotions based on self-labeled non-lexical vocalizations (Johnson et al., 2023).
Using earlier language modeling techniques, researchers would get their data from a variety of sources, including children’s stories (Mühlenbock and Lundälv, 2011), crowdsourced conversations (Vertanen and Ola Kristensson, 2011; Vertanen, 2017), online sources like Reddit (Yusufali et al., 2023), and TV shows (Yusufali et al., 2023). Recently, large language models (LLMs), such as ChatGPT, have been used for speeding up the writing process for AAC (Cai et al., 2022; Kumar et al., 2022; Valencia et al., 2023; Yusufali et al., 2023). AAC users are enthusiastic but cautious about employing LLMs to assist with their communication (Valencia et al., 2023). LLMs are extremely capable of generating coherent text, but there can still be issues of controlling the model to generate exactly what you want, especially if you are not used to using LLMs. Despite these advances, very little work has been done on non-English AAC (Amery et al., 2022; Mühlenbock and Lundälv, 2011).
There are also recent improvements in other aspects of AAC beyond the language component. The research in these areas include finding better interfaces (Gevarter et al., 2020; Light et al., 2019; Yusufali et al., 2023), improving the robustness of apps on smart devices (Laxmidas et al., 2021; Roark and Gutkin, 2022), and incorporating additional modalities such as physiological sensors (Holyfield and O, 2022), gestures (Ascari et al., 2021), or photographs (Fontana de Vargas and Moffatt, 2021; Pitt and McCarthy, 2021) and pictograms (Martínez-Santiago et al., 2016; Sampath et al., 2012) in addition to or instead of text.
3. Methods
We setup an in-depth interview to ask autistic adults how AAC applications can be improved. This resulted in six sets of questions. The questions were then verified by two autistic researchers who specialize in qualitative work who were not affiliated with the study. The full set of questions can be found in Appendix B.1, but the sets are summarized as follows:
  1. (1)
    Questions asking the participant to recall a recent conversation where they used AAC and describe details about how that conversation went.
  2. (2)
    Questions asking the user to think of particular scenarios when they used AAC.
  3. (3)
    Questions for describing a typical day of using AAC and what types of tools they used.
  4. (4)
    Questions asking about when they have felt empower or disempowered because of their AAC.
  5. (5)
    Questions about how they select what AAC to try.
  6. (6)
    Questions about privacy (ending with a privacy table).
The privacy table was a list of potential features that were presented to the participant after they had a chance to think about the privacy of their AAC on their own. For each of the seven features we came up with, the participant had to rate whether they would allow it in an AAC application they would use. More information about the table and its results are presented in Section 5.9.
We recruited participants via the Facebook group “Ask Me, I’m an AAC user!” and Twitter. The participants were asked to fill out an initial intake questionnaire (Appendix Section A), and if they qualified, they were emailed the consent form and a video of the first author reading the consent form aloud [^2] This practice of recording the consent form was taken with permission from Charlotte Brooks, which was well-received by autistic adult participants.. If they consented, they were asked to schedule an interview and were given a PDF containing all the interview questions & a PDF with the privacy table—the latter of which they were told not to open until the appropriate time. Participants were allowed to script their answers to the questions ahead of time, and were especially encouraged to do so if they had planned to use AAC during the interview.
We kept the recruiting open until we had 12 participants  [^3] It has been shown (Ando et al., 2014) that themes begin to stabilize after around 12 participants.. Seventy-eight people signed up for the study and filled out the initial intake questionnaire, twenty-five people were invited to participate in the interview, and twelve people were interviewed. For confidentiality, we have carefully revised the anecdotes to eliminate any information that could potentially identify the participants.
All participants were native English speakers, were living in the United States, were between the ages of 18-44 years, and used some form of high-tech AAC at least occasionally. Only two participants (2) said they used AAC full-time; the rest said that they were part-time users. A list of each participant’s age range, self-identified speaking label, current AAC, and amount of time using AAC (both in total and for the current AAC) can be found in Table 1.
Most of the interviews were performed via Zoom with video and audio. The interviews lasted about an hour, with the average being 1 hour and 18 minutes (standard deviation of 31.2 minutes, minimum 50 minutes, maximum 2 hours & 40 minutes). Interviewees were encouraged to use their AAC during the interview. One interview was conducted entirely asynchronously via email & text document. Another interview was conducted half on Zoom (cut off at 1 hour) and half asynchronously. For consistency, all interviews were run by the first author, who was careful to note the language the participant used when talking about how they referred to the use of mouth-words.
3.0.1. Analysis
We analyzed the interviews through thematic analysis (Braun and Clarke, 2006). We used Zoom’s built-in automated transcription tool to get an initial draft of the transcript. The transcripts were then corrected and verified by the authors. If an important section of the interview could not be confidently transcribed, the authors asked the interviewee for clarification via email later on. The authors independently formulated an initial set of themes by analyzing the transcripts in an asynchronous manner. Subsequently, they engaged in discussions to merge their respective themes, followed by a process of refining these combined themes and evaluating their alignment with the transcript content. Through iterative revisions and affinity diagramming, the authors reached a consensus on the final set of categories and themes presented in this paper.
4. Types of AAC Applications
There appears to be a divide between two categories of AAC applications based on input type: those that use typing-based input and those that use symbol-based input. Typing-based apps enable users to input any word they want on the fly, while symbol-based apps require pre-programming of what words will be displayed on buttons and what the layout of the buttons will be. The divide seems to revolve around two commonly-used applications (Proloquo2Go and Proloquo4Text), both developed by the company AssistiveWare [^4] https://www.assistiveware.com/. During the course of this research, AssistiveWare has come out with a new application in the Proloquo suite simply called Proloquo. Proloquo seems to solve some of the problems of this divide by beginning to merge the functionality of Proloquo2Go and Proloquo4Text. Proloquo is now a button-based layout similar to Proloquo2Go but without symbols (only words) and with more predictive text options. Proloquo4Text consists primarily of on-screen typing with the option to save phrases.
Other commonly-used applications among our participants include CoughDrop [^5] https://www.coughdrop.com/ by CoughDrop, Inc. and TD Snap [^6] https://us.tobiidynavox.com/pages/td-snap by Tobii Dynavox—which are both symbol-based apps. We have seen that those who prefer typing or swipe on phones & tablets usually find free applications that will support them in the ways they need. Please see Table 1 for the full list of applications and tools that our participants use to communicate.
Age Range Self-Assigned Label Years
Current AAC Years
18-24 Mostly non-speaking 5-9 Proloquo2Go,
Proloquo4Text,
TD Snap (for Swedish)
2-4
25-34 Unreliable speech since childhood writing,
ASL (working proficiency),
UbiDuo (device),
BuzzCards,
Relay
5-9
25-34 Semi-speaking 5-9 Proloquo4Text,
Proloquo2Go,
Tell Me
$\leq$ 1
25-35 Mostly non-speaking $\leq$ 1 Proloquo,
Proloquo4Text,
Relay,
CoughDrop
$\leq$ 1
25-34 Selectively mute/situationally mute 10+ texting/messaging,
Speech Assistant,
Proloquo2Go,
writing
2-4
25-34 Semi-speaking 10+ CoughDrop,
Predictable,
Speech Assistant,
many low-tech options
2-4
25-34 Semi-speaking $\leq$ 1 My Voice,
email,
Relay
$\leq$ 1
35-44 Mostly non speaking [with] unreliable speech and word retrieval issues 10+ Proloquo2Go,
TD Snap,
CoughDrop
2-4
35-44 Sometimes-user 5-9 Proloquo4Text,
Proloquo2Go,
Relay
2-4
35-44 Mostly non-speaking $\leq$ 1 ClaroCom,
Proloquo
$\leq$ 1
35-44 Mostly non-speaking 5-9 Proloquo2Go,
Proloquo4Text
5-9
35-44 Unreliable speech $\leq$ 1 TD Snap,
Speech Assistant
$\leq$ 1
Table 1. List of participants with their age range, self-assigned speaking label, years spent using AAC, their current AAC applications (the underlined app/technique being what they mostly use), and the number of years spent with their current application. Words in red are phone/tablet applications. ASL stands for American Sign Language. Relay is a free federal service for communicating with text over the phone.
5. Results
This section delves into the categories identified during the qualitative analysis of the transcripts. We will use a numerical value enclosed in parentheses is provided to indicate the number of participants supporting that particular theme (e.g., ”(3)” signifies ”three interviewees”). We will not be using any individual IDs due to the close-knit community of AAC users. We will also be referring to everyone by ”they/them”, even though we collected their pronouns during the interview. Direct quotes will be found inside quotation marks and italicized. The complete list of all 12 participants can be found in Table 1.
We have identified 9 main categories of themes, which we will describe in more detail in this section. The categories are: input options (§5.1), output options (§5.2), selecting or adapting AAC for a good fit (§5.3), when to start or swap AAC (§5.4), benefits (§5.5), access (§5.6), stumbling blocks for continued use (§5.7), social concerns (§5.8), and lack of control (§5.9).
5.1. Input Options
All 12 participants had preferences in terms of what input type they preferred or would like to see as an input option in a future AAC application. In this section, we discuss the use of keyboards, buttons/symbols, and other more indirect types of selection in applications, the differences felt between using symbol-based and text-based AAC, and aspects that make input easier such as vocabulary & layout flexibility, predictive text, or pre-programmed phrases.
5.1.1. Typing
All participants (12) wanted some type of typing input option to use at least occasionally, even if it is to augment their vocabulary for when they don’t have an appropriate button. Although a few (3) prefer a physical keyboard (“I can type on a phone, but I type 100 words a minute on a regular keyboard so that’s always gonna be my preference.”), most prefer or at least default to using an on-screen keyboard (8). One person (1) would take any kind of keyboard access “even if it’s an on-screen keyboard” but finds a physical keyboard better for their wrists and hands. Within those who prefer on-screen keyboards there was some variation in preferences of how that keyboard is implemented. Some (3) prefer a keyboard embedded within the application itself. One participant (1) said this is preferred because a native on-screen keyboard would mean “needing to swap motor plans back and forth a lot” (1) [^7] Motor planning is the process of learning how to perform an action or series of actions—or as one participant put it: “Your body knowing how to body.. A couple participants (2) had gotten around this by embedding a keyboard in the homepage of their AAC as regular buttons (for example, Figure 1). One participant (1) finds this so important that:
The biggest thing I looked for when I was initially selecting an app was the ability to add a keyboard to the homepage.
Figure 1. One participant showing their on-screen keyboard embedded on the homepage in Proloquo2Go.
[…]
A picture of an iPad with the program Proloquo2Go. The program has a big empty rectangle at the top for seeing the text that the user is currently writing to say. It is currently empty in this picture. The rest of the application is a giant grid of uniform buttons. There are categories of buttons at the very bottom of someone’s board. The home screen of their app has a lot of common words with stick figure symbols to correspond to the words. On the bottom of the grid (above the categories) there is an on-screen keyboard laid out in yellow buttons.
5.1.2. Symbol- vs text-based AAC
On the other hand, some participants have a preference for symbol-based AAC. One person (1) said that symbol-based apps work better for their brain but are much slower. Another person (1) said that they prefer having some images, although it’s not a requirement. A few participants (3) explicitly mentioned the difference they feel when using word- vs symbol-based AAC. They find typing faster (2) but compared to symbol-based buttons, they say typing creates “more mental effort” (1) or can be a “less calm” experience (1). One participant (1) will use symbol-based AAC around people because “there’s less pressure for me to find the correct words, and I have that extra support.
I like [button-based] Proloquo, though, because it forces me to slow down and think about what I’m trying to say. And for now, [typing-based] ClaroCom, when I have a thought already and need to get it out.
Some participants (7) spoke about supplementing or mixing symbol usage with typing, which can include things like using symbol-based AAC with Zoom’s chat (1) or completely mixing the two types of AAC (“Future AAC that has both symbols and typing and a vocabulary designed for autistic adults would be very empowering.”). Currently, having applications focus solely on symbol-based or text-based input slows users down.
[I dislike] symbols and typing being so completely separate in different apps. Having multiple apps is good and choice is good but sometimes its just like I don’t want to have to keep switching, and the typing in the symbol app just isn’t sufficient for what I need.
Others will actively avoid symbol-based AAC (2), although one participant (1) would accept an AAC that uses buttons without symbols. One participant said that since they see words in their head before they say them, it makes working with pictures more difficult.
5.1.3. Flexibility with vocabulary
One downside to using symbol-based AAC is that it could be limiting in the amount of vocabulary available. Seven participants (7) look for applications with high flexibility in terms of vocabulary, which may include an adult-level vocabulary (2), vocabulary for specific identity groups (1) or occupations (1), or an overall comprehensive lexicon (all 7 listed above). One participant (1) likes to use text-based AAC to see what they’re missing from their symbols. A couple of participants (2) have had experiences with older AAC with severely limited vocabulary, which they now know to avoid.
5.1.4. Layout
There were a lot of different preferences for how users want their buttons to be laid out. Some of these users wish existed in their preferred app or wish existed in general, others already exist and could be the reason why they chose the app that they use. Important aspects to consider for layouts include:
  • Grid size: The ability to set the number of buttons that appear on a single page (3) or pages that have different sized grids (1).
  • Organizing words: The order or groupings of what words appear near other words, such as alphabetically (1), semantically/associated (3), grammar-based (2), or organized manually/ color-coded (3). Two participants (2) appreciate a motor planning-based layout to help them know where words can be found. One participant (1) modified their layout to hide buttons that are used only in specific situations: “I think I keep it hidden cuz I don’t travel very much”.
  • Easy-to-find words: A layout that prevents the user from navigating folders or scrolling a lot to find a specific word (2).
  • Easy to edit: Multiple buttons can be changed or moved at the same time (2).
  • Button size: Either buttons that adapt to fill the page that they’re on (1) or buttons of different sizes on the same page (1).
  • Good core board: The app starts with a good layout that needs minimal adapting (2).
  • Consistent organization: Having the same layout across applications/devices (2).
5.1.5. Predictive text
Participants also enjoy input that aids them with drafting what they want to say, whether it be auto-complete/word suggestion (4), or helping them find the button they’re looking for via a search bar (2). Unfortunately, writing prediction is not available in symbol-based apps (1) nor when typing on a keyboard (1). A couple participants (2) have noticed that the predictive text is coming from their native keyboard rather than their app, although users who have Proloquo (1) see that it suggests related words next to the normal grid of words.
Usually these predictions are tailored to a user, which can be useful but also concerning, especially when it’s saving context-specific or misspelled words. One participant (1) suggests that future AAC could have selective forgetting of particular words.
I’ve got it programmed to save [words, and it picks up on things that are part of my] dialect like ‘wanna’ or ‘gimme’ […] I tell it to learn all those words, and it also learns anything I’m, say, telling a doctor. So, you know, it does save those words and phrases, and sometimes I misspell a word, and it’ll bring it up the wrong word to tell me […] But the ability to maybe tell it ‘forget this particular misspelling’ would be useful.
5.1.6. Pre-programed phrases
There are some features that can speed up communication. Participants (10) spoke about saving whole phrases and commonly-used messages ahead of time to be used for multiple situations.
I also like that with the phone I can pre-program phrases before I enter a situation, so if I’m gonna order, I don’t have to type it as I go, I can type it all in and hand it to them. Or I can have some emergency phrases about, you know, please call this person if I need help.
A similar feature is the quickfire reply (1), tapping on a word or phrase to immediately say the word without drafting. This can be useful for a quick yes/no response or for sentences that you know won’t change but are tiresome to type out each time, such as self introductions.
Note that quickfire replies are slightly different from speaking the words as the message is being drafted, which participants (2) disliked—one of them called it “infantilizing”.
I also hate apps that speak every word as you hit the button or speaks each sentence as you type it. It messes me up a lot and distracts me. I need to be able to turn that off so it only speaks the message when I’m finished.
Furthermore, when they do use pre-programmed messages, the person from the above quote prefers to have the option to edit it in the moment before saying it.
5.1.7. Indirect selection
The “default” input method for symbol-based AAC appears to be direct select. That is, tapping directly on the button of the word the user wishes to say. However, even in a group such autistic adults who are not necessarily using AAC due to a physical disability, some users still require or prefer an alternative mode of input. Sometimes participants will switch to an indirect select input type because a dynamic disability such as seizures (1) or hemiparesis (1) prevents them from easily using direct select.
Examples of indirect selection include:
  • Eye Gaze: Using one’s eyes to select words. Some participants are interested in or have tried using eye gaze as input: one participant (1) uses eye gaze regularly, another participant (1) would like to try it but doesn’t know how, and a third participant (1) has tried it but it gives them migraines.
  • Automatic Scanning: One button is highlighted at a time with the highlight automatically going through the buttons until the user selects one.
  • Step Scanning: A form of scanning with two switches. One button is highlighted at a time with the user pressing a switch to proceed to the next button. Another switch is hit to make a selection.
One participant (1) will use scanning or eye gaze when they have poor motor control. Their step scanning setup includes setting the “top half of the screen to select and the bottom half to advance to the next selection”. Another participant (1) wishes they could step scan with three switches instead of two: “forward, backward, and select, because I sometimes need to not be using direct select[,] and eyegaze and switch select [i.e., step scanning] both don’t work for me and it is exhausting to use a mouse”.
5.2. Output options
The two main themes that popped up in terms of output options were having the ability to show text to conversation partners and to have customizable text-to-speech (TTS)—also called speech synthesis—that fit the user’s identity. TTS & speech synthesis are the technical terms for the technology that speaks a selection of text aloud. This enables the user to share their thoughts with an entire group instead of being confined to showing a screen or piece of paper. We’ll go into these two themes (Showing text and Voice identity) in more detail in this section, but there were other interesting features that our participants mentioned either as something they have experienced and liked or would like to see in future AAC applications. These include:
  1. (1)
    A ”hold my thought” feature that saved what they were inputting so that could offer a quick reply to what was being said in the conversation and then return to what they were drafting (2).
  2. (2)
    Having the TTS begin reading out the start of a message while the remainder of it was being written (1).
  3. (3)
    Being able to quickly repeat (via TTS) what they had just said (1).
  4. (4)
    Teaching the TTS how to pronounce certain words correctly (1).
5.2.1. Showing text
Many participants (8) use the text display feature of their AAC app instead of, or in conjunction with, the TTS. This can stem from conversation participants unable to understand the synthesized voice and/or because the voice couldn’t be loud enough (6). Others prefer to show text to communicate most or all of the time (3) or switch between the different modes (1). The feature itself should also be flexible, such as with an easy adjustment of font size to show people from farther away (2) or the ability to flip the text around to show someone across from the AAC user (1). We discuss some of the limitations of showing text in Section 5.7.3.
5.2.2. Voice identity
A person’s identity is strongly tied to their voice. The availability of choice for voices (1) was an extremely important aspect for several participants, particularly not hidden behind a paywall or at least have the ability to be previewed before purchasing (1).
More customizable voices, for sure. It’s very personal and having the voice that matches every other person who uses AAC is very disempowering.
A couple participants (2) just want higher-quality voices, while others (1) know exactly what voice brands to look for (Acapela [^8] https://www.acapela-group.com/). One participant (1) has been using AAC for such a long time that they just want to use the same voice in order to sound consistent.
Although one participant (1) was happy that some voice packs were beginning to include a variety of different types of voices, one of the most common complaints we saw was that there aren’t enough nonbinary or middle-pitch voice options (4), and this becomes even more limiting when the app (such as TD Snap) limits basic customization such as adjusting the pitch (3). The only person who uses he/him pronouns out of our participants dislikes the men’s voices and prefers to pitch down child/teen TTS. Even if we were to sample their speaking voice to create a TTS voice, it might not match what they actually want to sound like:
I think the most important thing I look for is a compatible voice pack. As I mentioned before, I am trans, but I deal with age dysphoria as well, which is part of why I like [this voice]. Having a voice that sounds right is, therefore really, really important.
5.3. Selecting or adapting AAC for a good fit
Beyond input/output, participants have other criteria for selecting the AAC app that is right for them. In this category, we will go over themes around applications that fit the needs and preferences of the user.
Our participants described wanting to have a single or a few AAC application(s) that were flexible in design and would have the option to be used in multiple ways to suit their needs. This category describes themes where participants expressed the need for robust, adaptable features for their AAC so that they weren’t constrained to using an app that didn’t fit what they needed.
5.3.1. Aging up AAC
Participants (5) were of the opinion that AAC needs to be tailored for autistic adults, especially since they’re often targeted toward children.
Future AAC that has both symbols and typing and a vocabulary designed for autistic adults would be very empowering. Many AAC apps feel like they’re made for kids or students, and it feels infantilizing.
Two participants (2) described how the vocabulary available in the AAC tools had to be modified for adults. Another participant (1) felt that people who used symbol-based AAC as children might appreciate if there was a feature to switch over from images to text as they grow up. Someone (1) found that apps that report word frequency usage (“what words I need to practice or whatever”) was not a useful feature for adults who already know the language and was seen as another way of logging data. Another person (1) wished that there was more access to speech-language pathologists as an adult so that they might get more help for their AAC use.
5.3.2. Consolidation of features
Three participants (3) said that consolidating features from various apps into a single platform would make things more convenient, eliminating the need to constantly switch between different applications. Features are split across multiple applications. For instance, in one application, users could place buttons of different sizes on a single page, while another application would offer the ability to color-code items, which would be lacking in the first application. One of the participants (1) spoke about wanting to have all of the important features in one application.
There are things I like about each of the apps I use, but it would be nice to get all the features in one app. For instance, like I mentioned it’s a pain to use Proloquo2Go when you’re saving messages in advance, because you can’t easily see the whole message and everything later on. I also really wish that I could use Proloquo2Go on Swedish which I cannot. It would be nice to be able to have buttons in varying sizes on a single page, which you can do in TD Snap but not in Proloquo2Go and another feature that is missing from Proloquo2Go, which is in TD Snap is the chat feature, where if you’re in the middle of composing a message and someone interrupts that with another question, you can hit the chat button, compose a reply to the second thing, and then go back to composing the first message which had been saved where you left it.
There are also phone features that participants wished could be better integrated into their AAC applications to make their lives easier. These features included texting (1), phone calls (2), and the ability to easier play AAC audio from the phone speaker when headphones are being used (1).
5.3.3. General customization
All of our participants (12) expressed various levels of interest in overall customization and ease of use.
I like being able to adjust pretty much every feature really. Whether that’s the appearance of the buttons, how the grammar support behaves, how to navigate into and out of folders and so on.
Participants also expressed various preferences for how their ideal AAC application should look:
  • Preferred symbol library (4) - Boardmaker symbols, SymbolStix symbols (2), one with multiple skin tones (1), or having multiple library options—one participant (1) considers this when recommending an app to others.
  • Not overwhelming (7) - Interface that is not visually overwhelming (2), such as a clear layout and organization with fewer possibilities (4), or no bright colors (3)—especially when they are already feeling overwhelmed.
  • Easy-to-see symbols (2) - Symbols are high contrast (1). Another participant (1) avoids symbol-based AAC as someone who is low-vision.
  • Other visual customization (2) - Such as font selection (1) or dark mode for people with light sensitivity (1).
5.4. When to Start, Swap, or Stop AAC
Researchers still don’t know why autistic people may have trouble speaking orally, although some suggest that it could be anxiety-based (Muris and Ollendick, 2021; Ludlow et al., 2023)—although it’s hard to say whether anxiety is creating the difficulties or that having difficulties speaking orally gives people anxiety.
Melfsen et al. (2021)’s “unsafe world model” of selective mutism extends beyond anxiety and describes the lack of oral speech as a “shut down” of the nervous system from a perceived lack of safety. In this model, over-stimulation (among other things) can cause the perception of being unsafe. Related to this, two recurrent themes that came up in our study were the aspects of Trust of other people and the use of AAC to help them with Accommodating to current needs, which we will discuss in this section.
A third theory is that selective mutism could come from a potential weakness in neural auditory pathways (Muchnik et al., 2013; Henkin and Bar-Haim, 2015). In this section we will also discuss reasons why people might switch from one type of AAC to another—outside of input selection (§5.1) or output selection (§5.2).
5.4.1. Trust
Nearly all of our participants (11) spoke about how using AAC depended on whether the environment is safe, what the comfort level is with the person they are speaking to, and social safety.
When I don’t use AAC is in some situations when there are people who have dramatic power over me who can drastically control my life, because my trauma tells me it is actively not safe to in those situations.
As soon as that trust is broken, users would rather force speech than use AAC in front of them.
My friend said it felt a little incongruous to hear a child’s voice from [it]. So I only used AAC with them twice.
Some participants spoke about being more likely to use oral speech (3) or speak orally alongside AAC (3) with people they trust and in private situations. They are comfortable transitioning between speech and AAC in the presence of those who create a comfortable environment for communication.
Participants (2) don’t trust strangers to accept their AAC well; with good reason since they have been treated negatively in the past.
I didn’t use AAC last time I went to urgent care because I was worried the staff would treat me worse. […] I just used mouth words, and it went fine, but it felt like betraying my needs in a sense.
Other times it’s more nuanced. They feel pressured to be efficient in public (1) or are drained from matching others’ communication expectations (1).
Participants will often avoid social situations (4) rather than deal with strangers’ reactions to AAC. One participant (1) even avoids using AAC because they had abusive speech therapy when they were younger. This results in people not using AAC as much as they might need (2): “A lot of the time when I’m having to pull out AAC, I do the mental calculation of like ”do I really need to?””.
One participant (1), who mostly uses writing as their AAC, was the opposite. Because their speech is “unreliable”, they will primarily write with people who they are setting expectations with (1). Perhaps it is because of the lack of stigma against writing (vs an application) they are able to do this. Similarly, another participant (1) will either avoid social situations or use their notebook more.
5.4.2. Accommodating to current needs
There were themes of using energy to force speech or speak fluidly (4). Switching to AAC can feel like more (1) or less (1) energy than using mouth words. One participant (1) would occasionally fall into old habits of speaking orally by default. For another (1), who has an auditory processing disorder and mental fatigue, reading can be less tiring than listening. AAC can make it easier to overcome the “initiation cost” of figuring out when to talk (1).
Participants (9) mentioned losing the ability to use mouth-words because of various internal or external stressors. Some (2) spoke about using AAC when they were overwhelmed or had a build up of stressors from throughout their day. Others (2) spoke about their co-morbid conditions that trigger the use of AAC. Some participants (2) use AAC depending on their energy and effort required.
It’s important for users to have access to AAC tools and be able to identify when their bodies are reaching a tipping point so that they know when to start using AAC and select what AAC to use (2). One participant (1) will tell their partner to hold themselves accountable, making sure they’re respecting their body’s boundaries.
5.4.3. Accommodating to the environment
Several participants (9) spoke about how the environment such as the weather and noise affects their AAC usage. Participants (9) spoke about using their AAC tool to show text in situations where the environment is loud (1), in a drive-thru (1), when they preferred not to use text-to-speech and when they have been misunderstood. Some participants spoke about the limitations of displaying text via AAC which include sun glare (1), font size, and concerns about privacy when handing one’s phone to read the text.
Some examples of situations where our participants might not have access to their device include: services where body is in a particular position and a device can’t be reached or looked at (e.g., salon (1), eye doctor (1)), when their hands are full already (2), while walking (1), when there’s no place to put a keyboard/larger device (2), wet environments such as showers and pools (2), and cooking (1).
5.4.4. Accommodating to the task
A quarter of our participants mentioned that they had encountered problems using AAC to give a formal presentation (2) or prefer a certain application for presenting (1). One participant had mentioned that passing around a piece of paper for a presentation is not practical, especially in a group presentation. However, they ran into technical difficulties when the AAC device they had at the time (DeltaTalker) spoke the entire presentation too fast to be heard. Another participant (1) complained about how Proloquo2Go’s text field only lets you see a sentence at a time, making it hard to edit an entire talk. Similarly, someone else (1) will use Proloquo4Text when they wants to use more technical vocabulary. These examples highlight an issue that AAC applications make an assumption that someone is using them for short sentences and not longer forms of communication.
Other users will switch to alternate ways to communicate that are appropriate for the task, beyond specialized AAC applications. One participant (1) has shifted to having most important work conversations via email. Another participant (1) will use printed-out laminated cards of pictures of groceries to help themself make shopping lists.
5.4.5. Accommodating to conversation partners
Our participants would also switch AAC types depending on who they were communicating with. This need to switch can stem from accommodating the conversation partner’s disability, repairing if the partner cannot understand the TTS voice, or adapt to their partner’s skill of using a particular modality.
For instance, one participant (1) spoke about having options to adapt to others’ needs.
[Someone I know] is dyslexic and functionally illiterate. So needing to have both reading and auditory as options easily [available] and how to not have it overloading to switch between those for everyone based on what’s needed […] [It’s] a lot of: just what is needed in this moment? And having backups for backups and just being [ready] to switch in any moment.
Another participant (1) will switch to texting if their spouse doesn’t have their hearing aids in, or show them text if they don’t have their phone. A third participant (1) told of a story where they showed the text on their phone to a cashier, but she couldn’t see it and needed to take the phone to bring it close to her face. In order to communicate when their conversation partner can’t understand the TTS from their AAC, participants will change the speed of the TTS voice (1) or use chat/text to write out what was said (2).
AAC users will also change their conversation mode to adapt to someone’s skill level. One participant (1) is very comfortable with American Sign Language [^9] We recognize that ASL is a full language in its own right. However, it seems like our hearing participants use ASL more as a way of augmenting English., but they will write if people don’t sign and will practice signing if their conversation partner needs practice. Participants (2) also mentioned switching to whatever application their conversation partner is using in order to model/teach them how to use it.
5.5. Benefits
Of course people wouldn’t gravitate toward finding and using AAC if there weren’t significant benefits. The themes we found here include Organizing thoughts, Agency and identity, Self-accommodation, Community, and Fun.
5.5.1. Organizing thoughts
Sometimes I talk faster than I think and struggle to explain a complex idea, and AAC lets me collect my thoughts and write a clear message.
Participants (8) mentioned various ways in which AAC helped them “collect their thoughts”. The benefits can span from treating their AAC as an automatic note taking app (1) and keeping track of what they said (1) to getting around choice paralysis (1). A few participants will use their AAC to draft messages to be shared later via text or online (2).
Some autistic people have difficulty determining when it’s their turn in a conversation (Ochi et al., 2019). One participant (1) will use their AAC to type up what they would want to say, whether or not they will use it in a conversation. With having what they want to say out of the way, they can focus on when will be the right time to say it instead of keeping track of both what and when.
5.5.2. Agency and identity
Several participants (9) spoke about how AAC enabled them to communicate their needs and act independently. Using AAC has helped participants feel “like a whole person” (1) by letting them:
  • Be understood more (2)
  • Express their needs (3)
  • Decide what they want to say and how (1)
  • Speak for themselves and make their own choices (3)
  • Talk about heavy, emotional things (1)
Sometimes all it takes is having the ability to say something when it’s hard to otherwise (5): “Just still being able to contribute even on a bad day.
AAC can also become closely intertwined with a user’s identity:
Something which was really helpful to [trans people online was][…] being able to look at themselves using AAC as someone who is trying to deal with gender and how does that work when the world is neuronormative and assumes we speak as well?
5.5.3. Self-accommodation
Since communication is so core to our experience as humans, some of our participants have structured their entire lives around accommodating their AAC use. One participant (1) has made it so that the majority of their “heavy lifting conversations” are via email, and another participant (1) has structured their life around AAC such that they “don’t run into the edges very often anymore”.
More than half of our participants (7) rarely even use mouth-words anymore, except with people they are close to (2), around people who don’t support its use (3), when they can’t access it (2), or when they need to communicate in a certain way such as quickly or with emotion (1).
Participants feel more confident speaking orally when they know that they have AAC as a backup (2). One (1) makes sure that they has their AAC with them when they’re in public. Another (1) will bring their AAC along in case the day becomes overwhelming.
5.5.4. Community
AAC gives users the ability to feel socially connected to others (3). Our participants also reported making friends with other AAC users (2) or other autistic people since they don’t judge others for using AAC (1). Our participants also feel a need to give back to their communities. A couple advocate for AAC use (2) or are an autism self-advocate at local medical schools (1). Participants reported helping out other disabled groups as well. One participant (1) had transcribed for deaf teens at an event in the past, and another participant (1) who is for all intents and purposes deafblind themself, is training to be a deafblind counselor.
5.5.5. Fun
A couple of participants even spoke to the fun aspects of AAC, playing with it and talk to pets (1) or even having AAC as a special interest [^10] In the autism community, a special interest is a topic or hobby that deeply fascinates someone, and they try to learn all they can about it. (1).
5.6. Access
Due to autistic adults finding and accessing AAC later in life, there is an additional overhead that autistic children do not run into—either because they are still a part of the US education system and therefore can more easily access speech-language pathologists, or their parents and caregivers take over the details regarding things like price and setup.
5.6.1. Institutional support
Support from speech-language pathologists (SLPs) or other institutions such as government programs in place to help people who use AAC or who are disabled in general was mixed. Several participants (4) spoke about how speech therapists/speech-language pathologists were unhelpful or even abusive, hindering their ability to use AAC or find the right AAC for themselves.
In my experience, […] speech therapists were not helpful. […] There’s a lot they can do to acclimate folks to AAC better. Understanding what it means to have part-time speech or unreliable speech, how it shows up, and how to work with it instead of against it—so that you’re working on total communication, rather than just speech because technically they’re speech and language therapists.
On the other hand, SLPs can create opportunities for AAC users. One participant (1) had an SLP that let them try out an expensive AAC so that they could decide if it was right for them or not first.
While one participant had found SLPs unhelpful, they spoke a lot about benefiting from state-level support. They were able to borrow technologies over the years that helped obtain AAC without having to buy it themselves. However, one participant (1) had run into people at a state-level disability organization who were not “taught how to treat us as humans”.
5.6.2. Pricing and inaccessible operating systems
Some participants spoke about being prevented from getting the AAC that they thought would best fit their needs because it was only available on an operating system that they did not have access to. They either wanted an application that was supported by multiple operating systems (i.e., cross-platform) (1) or was not being constrained to one operating system (3). This second point came up often in reference to AssistiveWare’s Proloquo suite.
Actually existing at all on Android. I feel like that has to be said actually. I actually trialed both Proloquo2Go and Proloquo4Text and found neither of them worked for me, even beyond the fact that iPads don’t generally work for me.
These applications are only developed for iOS. Therefore, users would need to invest in an Apple product, which is typically more expensive than an Android tablet or phone. One participant (1) felt discouraged from ever bringing their iPad around for fear of breaking it because it was so expensive. Some participants mentioned having a dedicated tablet—typically an iPad—for AAC (4), while others are interested in getting one eventually (1).
Similarly, even more participants (10) spoke explicitly about how affordability played an important role in choosing an AAC application. Participants would gravitate toward applications that are free (4), particularly if they are getting a separate app for their phone (1). There are downfalls of using a free app though; namely, fewer features (1). This limiting of features is especially seen in the text-to-speech voice options (3). Even applications that cost money to use will have “premium” voices that cost extra.
It’s also pretty messed up that they make you spend 40 more dollars, and you don’t even know for sure you’re going to like the voices in the expansion.
Other times, participants will pay for expensive apps when there is a big sale (as AssistiveWare apparently does annually; 2), are occasionally gifted them (2), or just don’t use the AAC they think will help them because it’s too expensive (1). Only very rarely when they are concerned about price will they buy it themselves (1). One participant would just get whatever AAC their mom could afford when they were a child.
One participant (1) ran into a weird intersection between operating system incompatibility, price, and government assistance. They were interested in trying Speech Assistant application, but “that’s free on Android and \$20 on iOS, and I’m not paying \$20 for software that’s free on another platform just because I was able to get the State to buy me the platform most people prefer for AAC”.
5.6.3. Universal access
Some participants (3) spoke about the need for AAC to be “more accessible to more people”. AAC should be integrated into society (1) and autistic adults should have access to SLP professionals (1). Additionally, we heard themes about a need for increased awareness regarding available AAC options, with one participant discussing lesser-known features that are already available to people:
And then the other thing […] about AAC that people don’t know is that a lot of smartphones have built-in features that they don’t know are there, like Text Relay […] for being able to make your own calls, and we need to help folks learn about those resources. […] Like, the government funds that, and people just don’t know it’s a thing.
Participants (2) discussed being introduced to AAC much later in life than they would have liked. Even after knowing about AAC and using it some participants (2) expressed uncertainty as to whether an AAC app existed with the features that they’re interested in.
5.6.4. Setup effort
Once people get access to an AAC application, they spend a lot of time setting the app up and making boards to fit their needs (4). With so many complicated pieces, it’s difficult for users to remember where everything is and what it can do (3). It makes for a big overhead for starting AAC or moving to another AAC application (2). One participant (1) looks at other apps or the boards of other users to get help making their own boards, but they wish that there was a companion app for their computer where they can more easily edit boards.
I would prefer an end-user program on the computer– […] There’s an option to import boards, but where am I importing them from? […] It’s not transparent as to what format it’s in. […] And you can’t import a Proloquo board to Speech Assistant. I wish there was a standard.
5.7. Stumbling Blocks for Continued Use
Once our participants found and setup their selected AAC applications, there were still major issues that they would encounter. These issues might discourage users from using particular features of the apps or might make them disregard the application entirely.
5.7.1. Speed
Nearly all participants (11) brought up the issue of speed with AAC, wanting “to be able to communicate faster”(1) and keep up in conversations (3). Even patient conversation partners can sometimes forget what an AAC user might be responding to (1) or get distracted by others in the conversation (1). We could see an inconsistency in speed with how long the interviews themselves took. There was more variance in the length of the interviews for participants who used AAC (see Figure 2), which could depend on how comfortable the participant was with their AAC and/or whether they scripted their answers to the questions ahead of time.
Figure 2. A box plot depicting the difference between interview times for participants who primarily used AAC during the interview (right, orange) and those who did not (left, blue). This data does not include the one participant who interviewed completely asynchronously.
[…]
Speaking participants (5 people) took on average 1.045 hours (standard deviation of 0.05) for the interview, and non-speaking participants (6 people) took on average 1.5197 hours (SD 0.6).
This issue of speed can stem from troubles with locating button words on the board (6), but there also just seems to be an order for how fast someone can communicate based on what method they’re using. Oral speech seems to be faster than any AAC (1) and can speed up AAC use (1), but in the hierarchy of AAC, external keyboards are faster than on-screen keyboards (3), swipe keyboards are faster than typing letter by letter (1), and spelling things out letter by letter—either on an on-screen keyboard (1) or a letter board (1)—is the slowest. One participant (1) also found hand writing to be faster than phone typing.
Although on-screen typing/text-based AAC might be used more often than symbol-based AAC for complex words or faster communication (1; see Section 5.1 for more discussion about symbol vs text input), others wish that they could type faster (2). Having an onscreen keyboard option for spelling out words to look up where a button might be or to avoid having to look for a button (3) or using auto-complete/predictive text can help with faster button input (3). The important characteristic here being that the keyboard can be easily accessed without significantly disrupting app use (1) or motor planning (1). Issues with finding buttons are exacerbated for people who only use AAC sometimes.
“I would need to use it regularly enough to memorize the symbol locations. Proloquo’s search feature, at least, does highlight the specific buttons to press to get to the requested word, which I think is really handy.”
There are also extra issues that can slow down communication, such as an app slowly loading (1), not having the board completely setup (1), or having to switch apps in the middle of a conversation (1).
5.7.2. Voice quality
Most participants (10) expressed their concerns with text-to-speech (TTS). Seven participants (7) felt that the available TTS voices were often hard to understand or hard to hear in general. They shared instances—such as during interactions with doctors (1) or in meetings (1)—in which these voices led to misunderstandings. Two participants (2) attempt to get around misunderstandings by changing the speed of the voice to be slower. Other misunderstandings can come from their conversation partners who are not used to synthesized voices (1) or have auditory processing disorders (APDs) (1). One participant accommodates to their partner with an APD by using a voice that has similar characteristics to the dialect of the area.
A couple participants felt that the quality of the voices was simply too poor to use them (2).
I’ve had times where I tried to pull out the AAC in a group conversation […] and then people hear it and laugh because they think I’m making a joke. […] It’s just like no one really knows how to react to it, especially when you’re a part time user. And so you’re like ‘sorry, I showed up more disabled today than you know me as’.
A few participants (3) wished their AAC could be heard over the phone. Another participant (1) has been hung up on in the past because people thought they were a robot. Some participants (4) get around this by using TTY-Based [^11] https://www.fcc.gov/tty-relay or IP Relay [^12] https://www.fcc.gov/ip-relay, a free service where state-certified employees read messages transmitted via texting or the internet, respectively. This gives people who would rather communicate via text a human voice to “speak” with over the phone. As far as we know, there are no AAC applications that integrate with Relay. One participant (1) explicitly wished their AAC could.
One participant (1) spoke about how the TTS lacks emotion, while another used a child’s voice because it was more expressive (1). Additionally, the voices from the device can get masked by background noise because they simply can’t become loud enough (4). Participants who primarily show text consider TTS as a bonus (1).
Figure 3. A participant showing sComm’s UbiDuo device. These devices can be sold in pairs with either a wired or wireless connection. Despite its bulk, they prefers something like this over a phone app because of their ability to type quickly.
[…]
The UbiDuo device has a slightly smaller than normal sized keyboard with a small (3-5 inches wide) backlit screen above it. It looks like the screen can fold down toward the keyboard, but whether it’s folded or not, the base looks like it might be at least an inch thick.
5.7.3. Limitations of showing text
Despite showing text being commonly used and even used as a back-up for TTS-based output, there are some severe limitations to it that auditory communication does not have. Screens are often too small to show text (1), which makes it particularly difficult to share with people at a distance (1) or sharing something with a group of people (2). Because of this, especially with phones, users have to hand their device to another person (2), which can be a security risk. Reading from a smart device can have extra complications since there can frequently be a glare on the screen, especially when outside (1).
5.7.4. Device restrictions
There are times when our participants (6) struggled with the physical devices themselves. This is becomes obviously when an AAC application is put onto a regular phone (vs a tablet). A phone’s screen is typically too small to see the symbols (2), but even using text-based AAC requires the user to scroll more to see text output (1), which is difficult when showing text to someone. One participant (1) wears headphones a lot, which are connected to their phone, and so it’s not easy to switch the audio to let others hear the synthesized voice from an AAC app. Another participant (1) finds the screen so limiting that they wish there was a computer program for their app so that they could edit boards on their PC before putting them on their phone.
One of our participants (1) has a particularly interesting device called the UbiDuo—shown in Figure 3, which has a standard keyboard and is not connected to the internet. However, it is incredibly bulky and single-purpose, unlike a tablet or phone.
I mentioned it was bulky, which created a lot of barriers. Number one, the weight, and then not everything fit in my backpack, I couldn’t have it on my desk and do my work. So, I’d have to get it up and down to use it.
Even tablets aren’t perfect as older devices can have unpredictable quirks (1), such as crashing in extreme temperatures (1).
5.7.5. Unreliability
One point that was repeatedly brought up was the wish for continued, reliable use for AAC (6). It’s so important for one participant (1) that: “I always have and always will carry pen and paper just because it’s the most reliable”. It’s especially important when even commonly-used applications have bugs. Several participants (5) ran into bugs such as sound glitches (1), external notifications breaking off TTS (1), or erasing what the user was typing (1). Although they didn’t mention bugs explicitly, one participant’s (1) device couldn’t say ”TTY” during the interview and they had to type it in the chat. It can be disheartening, as well, when bugs are reported but the request to fix them is dismissed.
I am by far not the only adult autistic AAC user who feels like Coughdrop support kinda blows us off. Like, in the three months I used Coughdrop, I reported several bugs and *never* got any fixed, it was always ”we’ll forward it but no promises when he’ll have time to look at it”.
5.7.6. Language choices
Most AAC is English-specific (Amery et al., 2022). People who speak multiple languages or are learning another language are stuck having to find another AAC application to speak in that language (Swedish; 1), attempting to make their own (Finnish; 1), or never talking in that language (Irish; 1).
In particular I’ve put about two years into learning Finnish but still can’t get any AAC apps in Finnish because [it’s] not like there’s many people who want to deal with any non Indo-European languages, but I know so many people who’ve run into this—including with Indo-European languages that are not popular languages.
5.7.7. External support
Customer support can be key to whether or not a user will keep using an application. Participants who experienced good customer support (2) feel like they can rely on the company to continually improve the app. On the other hand, participants who experienced bad customer support (3) will switch applications, if possible, or suffer in silence since their concerns are not being actively addressed by the developers.
Beyond the company themself, users might be able to find support through tutorials online (1), but less commonly-used applications have fewer users and therefore less of a community to draw from (1).
I wish there was […] an export feature where […] there could be a forum, so that I could share my board, and you could share your board, and everybody could.
5.7.8. Restrictive features
Restrictive features (1) or attributes of an app that make it harder to customize (2) can be significant hindrances. Occasionally, these can be okay if they can be turned off or unused without hurting the interaction with the app (1). A couple examples of restrictive features that people reported include having people unable to call someone back so that they can use Relay (1) and having to use the TTS for pre-programmed phrases because they don’t show up in the text box when selected (1).
5.8. Social Concerns
Many social issues arose even when prompting participants about how AAC technology can be improved. All of these social concerns can be boiled down to a lack of respect from conversation partners and other people in the user’s environment. Most of the time this lack of respect stems from ignorance. The majority of our society is not used to seeing people use AAC and therefore do not know how they are supposed to react. Many times if people had just acted normally it would have been better than doing what they believe is “accommodating”.
5.8.1. Communication not being taken seriously
Many participants (8) spoke about situations where they were not taken seriously by conversation partners. A few participants (4) have run into others treating them as if using AAC means they’re less capable of making decisions or being competent. In particular, there is a perceived stigma and infantilization around using symbol-based AAC (2)—“people seem to think symbols means you’re less capable or something”.
Meanwhile, others were concerned they wouldn’t be taken seriously had they used their mouth-words (1) or had switched back and forth between using AAC (1). For example, one participant (1) had used AAC in the past to make sure they made a good initial impression on their classmates and instructor: “I wanted them to take me seriously, and I knew that my voice would not come out clean and fluent and that would be held against me”.
5.8.2. AAC use not normalized
All of our participants (12) spoke about how AAC use is not normalized in society and described social situations where they encountered people not accepting or expecting AAC as communication. Participants will sometimes have to explain or have others explain for them that they’re switching to AAC (2) or will avoid using AAC because they don’t want to deal with explaining it to every person they run into (1). Occasionally people don’t understand why AAC is being used (2) or at least don’t expect it (2). Many people do not know how to communicate with AAC users and will assume someone who cannot speak orally is deaf or hard-of-hearing (4). Conversation partners will also, subconsciously or consciously, gesture more often (1) or use fewer mouthwords themselves (1) when speaking to an AAC user.
In the worst cases, conversation partners will not be respectful (4) or will outright not accept its use (3). Participants who occasionally use mouth-words will get weird responses when they need to switch to AAC at some point (3). Therefore, participants think there needs to be more education around how to respect AAC users (4) — “there are some cases where I need to train people to respect AAC, and I’m not sure about how to manage that yet”.
5.8.3. Prevented from using AAC
Due to the lack of normalization, several instances came up where participants were prevented from speaking because their AAC options were taken away (6), such as their whole device being taken (4). Two participants had their devices confiscated at hospitals because the device either had a camera in it (1) or because it was a “no phone space” (1).
One participant (1) had gotten around this in a courtroom by using low-tech letterboards, but even low tech options are not immune to being removed as an option. Another participant (1) spoke about how breaking their right arm in 7th grade—where writing was their primary mode of communication—meant an entire 6 months where they could barely communicate.
Occasionally devices are taken away so that others could read what they had written (1) or to even play around with the application (1). One participan t(1), who uses the bulky UbiDuo, recounted a harrowing ordeal at a basketball game they went to:
They search any bags you bring, and you have to walk a long ways to get there, so it didn’t make sense to try and bring a device outside of a phone. […] A UbiDuo would be too big and then they might freak out cuz it’s technology. So I went to this event, and it ended up being really loud. They had a loud halftime show, and I lost speech and a lot of motor skills. I needed a different way to communicate, and I couldn’t because I didn’t have the right tools. My hands would not cooperate for typing on a phone. My body was jerking around enough that they couldn’t read my signs.
5.8.4. Impatient conversation partners
Many participants struggle with impatient conversation partners (7) or have trouble getting people to listen/pay attention (4). AAC users will often have others talking over them (2), especially in groups (4). Occasionally people will try to guess what the AAC user is about to say in order to speed things up (1), but this will often make the conversation slower.
I remember going to meetings of the disability group I started in my school, and they just all kept talking so quickly and changing topics that it was frustrating to try to participate with AAC. And I felt like I either had to give in and use mouth words, or I wouldn’t be able to say what I was thinking”.
5.9. Control of Communication
Wanting to be autonomous and in control of their own communication, autistic adults have concerns about people and companies potentially having access to everything they say. Many AAC applications are made with SLP intervention in mind, as children are learning to speak for the first time with their AAC. However, adults finding AAC later in life do not need these supports and the conventional logging practices can seem invasive at best. We begin this section with a discussion about being in Control of features, where we will also dive into the privacy ratings we collected at the end of the interview and cross-examine them with our qualitative analysis. Unwanted outsider access and Company trust also arose as themes in the context of privacy.
5.9.1. Control of features
In addition to the semi-structured questions, we listed features that loosely aligned with modern natural language processing techniques and asked participants whether they would use them, with the framing of privacy (never, only in the current conversation, across conversations, or always okay). The features were: automatic speech recognition, logging conversations verbatim, logging the topic of the conversation (i.e., topic modeling), saving personalized information automatically, saving personalized information explicitly, using a history of sentences said in the past (i.e., language modeling), and using a history of words said in the past (i.e., bag of words). The full table that the participants saw can be found in Appendix B.3.
The results of the responses, in aggregate, are shown in Figure 4. Overall, deliberate personalization was the most accepted, which makes sense since a lot of our participants take the time to personalize their AAC with words that are unique to them. This type of information seems to be mostly okay for applications to store. Although the question of whether it’s stored locally or in the cloud might change their use.
Figure 4. Level of comfort for use of and data storage from 7 potential features. Participants rated each feature as either something they would never use (red), something they would only use as long as the data persisted only in their current conversation (orange), something they would use with the data persisting across conversations — with the assumption that it would be deleted over time (yellow), or something they would be okay using all the time (green). The specific descriptions for each feature can be found in Appendix B.3.
[…]
Number of people who would use each feature. Automatic personalization: 6 never allow, 2 current conversation, 3 across conversations, 1 okay Speech recognition: 5 never allow, 4 current conversation, 1 across conversations, 2 okay Log verbatim: 2 never allow, 4 current conversation, 6 across conversations, 0 okay Log topics: 2 never allow, 1 current conversation, 8 across conversations, 1 okay Sentence prediction: 1 never allow, 2 current conversation, 5 across conversations, 4 okay Word prediction: 1 never allow, 0 current conversation, 5 across conversations, 6 okay Deliberate personalization: 1 never allow, 0 current conversation, 2 across conversations, 9 okay
While talking about privacy (starting at Question 6), some participants (4) mentioned how they were okay using features as long as they’re explicitly used and can be turned off easily. People don’t want to be forced into certain features (1) and they want the ability to have control over details of how a feature is used (1).
I don’t like it when the way I use something automatically changes the way an app or website, or whatever, functions. I’m fine with it if I’m choosing to make the changes like deliberately adjusting settings or changing the layout of AAC app, for instance, but if it’s automatically adjusting itself in response to how I interact with it, I hate that.
Further, automatic personalization was the least liked in our quantitative analysis (Figure 4). Half of our participants said that this was never a good feature, more than any other feature.
Some participants would even occasionally allow invasive data collecting such as the logging of conversations as long as they had control over what it was logging. For example, if there was a way to delete items (1) or to make sure certain conversations aren’t logged (1). Others would never allow logging that couldn’t be turned off (1). Still others were fine with logging as long as it didn’t analyze their data (1). However, none of our participants were comfortable having logging verbatim always on (Figure 4).
I turned off every single prediction that ClaroCom, including its built-in support for learning automatically. I don’t even like that it has those knobs. I’d rather it just plain didn’t do any of that. But unfortunately, needs must.
Speech recognition also had 5 ”never” ratings, perhaps because of the thought of the app “listening in” on other conversations or on conversation partners who did not consent to be transcribed.
5.9.2. Unwanted outsider access
Although some participants find company data collecting useful, participants do not want this collection to happen automatically. Most participants (11) were concerned about outsiders either getting a hold of their data without their consent. Most of the time this concern is centered around companies logging data. While data security was certainly a concern, participants expressed explicit concerns about logs of conversations ending up in the wrong hands—such as authorities (1), parents (1), or “others” such as SLPs (1). The concern being that others using these logs can use them to abuse the AAC user (1).
The concerns were raised about sensitive data being saved. One participant (1) wanted to make sure that they had legal access to the things that they say. Another participant (1) was less concerned about their own words being seen and more concerned about the information they share with other people. At their job, they had to take extra measures to make sure that their AAC was HIPAA-compliant:
I do have to think a little about data security specifically for my job because I’m bound by HIPAA and some stuff I say is confidential, but the HIPAA compliance person at my job, talked through that with me to make sure that things were in compliance.
People (2) were also concerned about tracking data that they felt was unnecessary for an AAC app to track, such as location. Other participants showed concern about people reading what they’re saying before they say it (2), other parties reading text intended for conversation partners (1), or conversation partners going through their phone when they need to show them the screen or share their device with them (2):
Sometimes my AAC use involves handing my phone to another person for them to read and I always have to trust that they’re not going to then use my unlocked phone and go snooping.
Only one participant didn’t express immediate concern about outsider access. They were the only participant who has been using various types of AAC since childhood. They have used a variety of AAC technologies their whole life, most of which were created pre-internet and only had data stored on the device. Now, their AAC use involves writing on paper and using services such as Relay that have professionals reading their text aloud.
Well, they’re held a really strict confidentiality standards, so I never have concerns about that. […] [they] can lose [their] license […] And if the sensitive information I would be sharing on Relay would be medical stuff, it’s a stranger.
5.9.3. Company trust
Some participants (4) trust the company of their AAC applications well enough. One participant (1) can tell that the company of the app they use (AssistiveWare) appreciates feedback, and this helps them know that they care about their user base. However, even well-intended companies might slip up sometimes. One participant (1) told a cautionary tale of a company misusing data. That is, they thought they could use anonymized data without permission (1).
Many tech companies rely on data to train machine learning models, but even without training models, companies need to know how people are using their applications and how they can be improved. So what type of data are AAC applications saving about their users? Nine participants (9) mentioned some uncertainty about the data policy of their AAC application, either in general or in the specifics of particular features of the app. So it makes sense that people (2) would appreciate clear data policies. One participant (1) teaches people how to clear the cache for their AAC apps so that they can have some control over what their phone is storing.
6. Guidelines for Improving AAC Use
In this section, we discuss a set of recommendations based on our analysis for how AAC features can be improved from a research perspective, how AAC should be designed and implemented for autistic adults, and what policy makers might be able to do to improve access and normalize AAC use.
6.1. Guidelines for Natural Language Processing (NLP) researchers
As large language models (LLMs) like ChatGPT become more prevalent in AAC technologies, we need to be able to integrate these new methods while also keeping the user centered. In terms of what autistic adults need from the NLP research community, we have identified three major thrusts.
6.1.1. Faster communication
Although there has been a lot of work in word prediction (see Section 2), there is still a long way to go to integrate them smoothly into AAC applications so that there is no additional overhead. Even though LLMs can produce much more coherent text, they take a lot of compute power and time to run. Therefore, despite the potential for using LLMs in AAC, they are not a viable option when it comes to speed and the ability to run on a phone. If LLMs are going to be used, they must not be during a live conversation. Smaller models will have to be used for on-the-fly word prediction.
6.1.2. Transparent data collection
Machine learning models need data to train on, but nobody wants every single conversation that they’ve ever had to be recorded and put into a model for everyone to use. Therefore we need transparent data collect where we tell users exactly what will be recorded if they opt-in to data collection in the first place. It should be explained in as much detail as possible (and with no unexplained technical jargon) what the data could potentially be used for. Furthermore, there should be an obvious way for the user to temporarily not collect data and/or remove certain sentences from what was collected. Certain sensitive topics might end up with sparse data, but perhaps users can put up with less predictive for those specific conversations — as long as it is explained to them.
6.1.3. Voice customization
AAC devices should provide inclusive and diverse text-to-speech voices so that users can find a voice that matches their identity. They should have the ability to use a voice that is unique to them and can be used across multiple devices. The voices should be easy to understand, able to be played at a loud volume, and a realistic as possible.
6.2. Guidelines for AAC application makers
Here we will provide suggestions for anyone who wants to make AAC applications, regardless of whether you’re a company, a researcher, or an individual developer who wants to make their own application.
6.2.1. Transparency of features and data
First and foremost, it is essential for applications to be transparent with what data they save, what data they store locally on the user’s device, what data they store on a separate server, and if/how the data is being processed. All of this must be described in plain language (no tech/legal jargon) so that users can understand what to expect from the application and to build trust in the company who made the app. No data should be collected without the explicit permission of the user. Beyond this, any feature that is not considered part of the core functionality of the application must be “opt-in” rather than “opt-out”. That is, users must know that the feature is there so that they can decide whether or not to turn it on.
6.2.2. Modular design
Since all of our participants valued high levels of customization and a variety of features to fit their needs, we recommend that application makers adopt a modular design. Although the implementation is non-trivial, modular AAC applications would enable users to customize the app to their liking without sacrificing the needs of other users. This can be done in part by allowing for third-party developers to create plug-ins. These plug-ins would have to follow the transparency guidelines outlined in Section 6.2.1. This will encourage users to augment their AAC and share their modifications with the community. It will create standards across users so that even people with very different apps can share components they find useful without a large overhead.
6.2.3. Adult-first design
Many existing AAC applications are created with children in mind, and this leaves autistic adults with having to force applications to fit their needs. We suggest that there should be some AAC applications that are designed for adults first, since they are just as important of a group of users as children are. Furthermore, children will eventually grow up and will most likely want to shift to AAC that reflects their development. Design decisions like this might include less recording of button presses and less reporting functionality.
6.2.4. Variety
We would like to encourage more companies to create AAC applications. By having a variety of applications—in conjunction with making AAC use more commonplace, we can cover a larger suite of needs and drive down the price of individual app purchases. Once prices are cheaper, more people might be willing to pay for AAC.
6.3. Guidelines for policy makers
6.3.1. Greater support of use
We can support adult AAC users by providing more access to SLPs and other professionals who can setup AAC. These professionals should be educated on autistic adult perspectives so that they are informed before transitioning from child to adult patients or from allistic (not autistic) to autistic patients. There should also be a greater effort to make dedicated AAC devices free or affordable.
6.3.2. Broadcasting of existing accessibility programs
Participants spoke about situations where they weren’t aware of a certain feature/device being available to them. One participant highlighted that numerous users are unaware that smartphones often come with built-in features like Text Relay. Moreover, many AAC users are not informed about state government funds that they could access. It is essential to provide resources to educate AAC users or potential AAC users about the range of options already available to them and have a compilation of resources available depending on what state they live in.
7. Limitations & Future Work
Although we were able to collect some variation in gender and age, all of our participants were white. This is a significant limitation to our work, which we attribute to our recruitment methods. Therefore, any insight toward AAC shared in this paper may not apply to BIPOC users. This is particularly important since some (Pope et al., 2022) have identified racial disparities in access to such communication tools.
Another limitation is that our participants were only recruited online. As we did the majority of our interviews during the pandemic, there was not much of a choice. However, because of this we are missing out on a fraction of the autistic adult population that does not use the internet or social media. This could be why our participants skew on the younger side.
We also recognize that only 12 people were interviewed in this study. While we were able to get an in-depth look into some participants’ use of AAC, this only provides a snapshot of the types of design decisions we might want to consider when creating AAC applications. A much larger, shallower study must be done in order to understand a variety of perspectives.
We plan to create a prototype considering many of the features that arose from this study to see how people actually like these components in a potential AAC system and how they can best be implemented. We will also be sure to follow our suggested guidelines to determine if they indeed work and come up with new guidelines as needed.
We would also like to encourage the rest of the AAC research community to take our insights and guidelines into consideration so that we can iterate on them as a community. This needs to include support from existing and future companies that make AAC so that communication can be improved specifically for autistic adults and not as an afterthought.
8. Conclusion
Our interviews with autistic adult AAC users revealed both current societal and technical issues encountered with AAC applications. We identified what types of input & output options people prefer for their AAC, what considerations people take into account for when selecting an AAC application or switching to a different AAC application, the benefit of using AAC, issues with accessing or using AAC, social concerns, and concerns about control over their communication. We conclude by suggesting nine guidelines for natural language processing researchers, AAC application makers, and policy makers to make AAC better suited specifically for autistic adults.
Acknowledgements
This work was funded in part by the National Science Foundation under Grant #2030859 to the Computing Research Association for the CIFellows Project. We would like to extend a special thanks to Dr. Layne Jackson Hubbard and Dr. Alyssa Zisk for looking over the interview questions; Dr. Danaë Metaxa & Dr. Andrew Head for HCI methods help; and Dr. Chris Callison-Burch for being a great postdoc/masters mentor. Extra special thanks to all of our participants! We hope we did your words justice.
Appendix A Pre-Interview Screening Questionnaire
The following is the text shown to potential participants via a Qualtrics survey. Any text in this section that uses this font was not displayed to the user and is only included in this paper for additional context.
Hello! Researchers at the University of Pennsylvania are currently looking for autistic people to be interviewed about their AAC (Augmentative and Alternative Communication) usage. Please note that this study is on digital AAC applications, such as those used on phones, tablets, or other devices. If you are interested in being interviewed for our study, please fill out this preliminary survey to see if you qualify. Thank you!
  1. (1)
    What is your email address? If you qualify for the interview, we will contact you via email. [text entry]
  2. (2)
    Would you like to participate in our study?
    1. (a)
      Yes
    2. (b)
      No
  3. (3)
    Are you autistic? (self-diagnosed is fine)
    1. (a)
      Yes
    2. (b)
      No
    3. (c)
      Maybe
  4. (4)
    What is your age?
    1. (a)
      Less than 18 years old
    2. (b)
      18-24 years
    3. (c)
      25-34 years
    4. (d)
      35-44 years
    5. (e)
      45-54 years
    6. (f)
      55-64 years
    7. (g)
      65-74 years
    8. (h)
      75 years or older
  5. (5)
    Which country do you currently live in? [Drop-down list of countries]
  6. (6)
    Are you a native English speaker? (Have you been speaking English since childhood?)
    1. (a)
      Yes
    2. (b)
      No
  7. (7)
    Are you non-speaking (i.e., do not speak verbally) or lose speech occasionally?
    1. (a)
      Non-speaking
    2. (b)
      Mostly non-speaking
    3. (c)
      Selectively mute/situationally mute
    4. (d)
      Semi-speaking
    5. (e)
      Unreliable speech
    6. (f)
      Other AAC user [with free-form text entry]
    7. (g)
      None of the above
  8. (8)
    Do you use an AAC application to communicate part-time or full-time?
    1. (a)
      No
    2. (b)
      Part-time
    3. (c)
      Full-time
If the participant selected ‘‘Part-time’’ or ‘‘Full-time’’, they would be shown the following questions:
  1. (9)
    Approximately how long have you been using AAC for?
    1. (a)
      Since childhood
    2. (b)
      10+ years
    3. (c)
      5-9 years
    4. (d)
      2-4 years
    5. (e)
      1 year or less
  2. (10)
    How long have you been using your current AAC app?
    1. (a)
      Since childhood
    2. (b)
      10+ years
    3. (c)
      5-9 years
    4. (d)
      2-4 years
    5. (e)
      1 year or less
  3. (11)
    What do you look for in an AAC application in general? (There are no right or wrong answers!) [text entry]
  4. (12)
    Do you have any questions or comments for the researchers? [text entry]
As we were focusing on autistic adult AAC users who spoke English, any participants who reported ‘‘No’’ for question 2, ‘‘No’’ for question 3, ‘‘Less than 18 years old’’ for question 4, ‘‘No’’ for question 6, or ‘‘No’’ for question 8 were given an automated rejection notice. The rest of the participants were shown the message:
Thank you for your interest in our study! We will reach out to you shortly to schedule an interview. Please expect an email from [redacted email address].
Appendix B Interview Questions
Below are the questions that the participants were asked. Section B.1 shows the entire set of interview questions verbatim, including text formatting such as bolded text, underlining, and a footnote.
Mostly speaking and mostly nonspeaking participants were provided the same questions with the exception of Question 1. Section B.2 provides the alternative Question 1 that we asked participants who were primarily nonspeaking.
Once the participants had looked at all of the questions, they were allowed to fill out the privacy table (Section B.3) that was given in a separate PDF called ‘wait_to_open.pdf’.
B.1. Interview Questions (with Question 1 for mostly speaking participants)
For any question, if something doesn’t apply to you, please explain.
All of the questions relate to dedicated AAC applications and related tools you might use (such as word processors).
  1. (1)
    Think of a recent time when you used AAC.
    1. (a)
      What triggered your need to use it (if you’re comfortable with sharing)?
    2. (b)
      How did the conversation go? Were there misunderstandings?
    3. (c)
      Approximately how long did you use it in the single conversation?
    4. (d)
      If you speak orally as well, what are ways AAC has supplemented your conversations? i. Are you using AAC alongside or instead of speech?
    5. (e)
      What other ways did you communicate (e.g., non-verbal communication)?
  2. (2)
    The following questions will ask you to think of specific situations.
    1. (a)
      Tell me about a situation where you wanted to use AAC but couldn’t because of the limitations of the application(s).
    2. (b)
      Tell me about a situation where you wanted to use AAC but couldn’t because of the type of social situation.
    3. (c)
      Do you use AAC in private, out in public, or both?
      i. If you haven’t already, tell me about a time you used it in private.
      ii. If you haven’t already, tell me about a time you used it in public.
      iii. How did these two times differ? Was one easier than the other?
    4. (d)
      Tell me about a time when you used AAC but it didn’t work.
      i. For example, the AAC application was not sufficient for your communication needs or even the application itself was broken in some way.
      ii. Describe some things that you dislike(d) about the app/tool.
      iii. Have you continued to use the app/tool? Why or why not?
  3. (3)
    Describe a typical day when you use AAC.
    [This is the part where I’d like pictures showing me your application(s). It can be a screencap or a picture taken from another device.]
    1. (a)
      What app(s) do you use?
    2. (b)
      If you have multiple AAC apps, how do you decide which one to use?
    3. (c)
      Can you show me or tell me about what you like about your app(s)?
    4. (d)
      Is there anything that you wish the app could do that it doesn’t do?
    5. (e)
      Do you think this AAC app limits you in some way? If so, how?
  4. (4)
    Empowerment 111“the process of becoming stronger and more confident, especially in controlling one’s life and claiming one’s rights” – Oxford Dictionary
    1. (a)
      Tell me about a time AAC helped you feel empowered as an autistic person.
    2. (b)
      Tell me about a time AAC made you feel disempowered as an autistic person.
      i. Now imagine a brand new AAC tool that could have helped you in that moment. What would it be like?
    3. (c)
      How do you imagine future AAC could empower you as an autistic person?
  5. (5)
    Think about the types of AAC applications you have used (both now and in the past).
    1. (a)
      What types of features do you look for in an AAC app or tool?
    2. (b)
      What features do you try to avoid?
  6. (6)
    Have you ever been concerned about your privacy when using AAC?
    1. (a)
      What type of information about you do you think your AAC application(s) saves?
    2. (b)
      Where do you think it’s stored (e.g., on your device, in “the cloud”)?
      i. How secure do you think it is?
      ii. How long do you think it’s stored for?
    3. (c)
      Is there any feature that you would never allow to be in an AAC application you use? Any that you do allow but would prefer not to have?
    4. (d)
      Once you answer Question 6c, open the file “wait_to_open.pdf” and rate how you feel about the features in the table.
  7. (7)
    Is there anything else you think I should know that we haven’t spoken about?
B.2. Alternate Question 1 for Mostly Nonspeaking Participants
  1. (1)
    Usage.
    1. (a)
      When do you not use AAC?
    2. (b)
      Think of a recent conversation you’ve had. Were there misunderstandings because of AAC?
    3. (c)
      If you speak orally as well, what are ways AAC has supplemented your conversations?
      i. Are you using AAC alongside or instead of speech?
    4. (d)
      What other ways did you communicate (e.g., non-verbal communication)?
# AAC Feature Never
Allow
Current
Conversation
Across
Conversations
Okay
1 using speech-to-text (i.e., having the app “listen” to people around you)
2 keeping a log of what has been said in the conversation word for word
3 keeping a log of the topic(s) of the conversation
4 saving personalized information (e.g. favorite TV show, best friend’s name) automatically
5 saving personalized information only when you explicitly permit it
6 using sentences you’ve said in the past to predict what you want to say
7 using words you’ve said in the past to predict what you want to say
B.3. Privacy Table
For each row of the table, read the feature and think about how you would feel using AAC with that feature—think specifically in terms of your privacy. The options are:
  • Never Allow - This is a feature you would never permit in AAC. It might even discourage you from choosing/using the app.
  • Current Conversation - This is a feature that you’re okay with as long as the data is used in the current conversation only and then deleted.
  • Across Conversations - This is a feature that you’re okay with across several conversations with the assumption that the data will be deleted at some point in the future—for example, either by you or by the app after a certain period of time.
  • Okay - This is a feature that you’re always okay with, regardless of the circumstances.
Put an X in the cell that best matches your preference for each feature. Feel free to elaborate on why or how you made your choices.